# Action2Sound: Ambient-Aware Generation of Action Sounds from Egocentric Videos

### v2

@@ 5.1 Evaluation
    Audio-visual synchronization (AV-Sync) [38]: a binary classification model that classifies whether the video and generated audio streams are synchronized. Following [38], we create negative examples by either shift audio temporally or sample audio from a different video clip. See more details in Supp.
@@
Neither learning-based model has the ability to tackle the ambient sound, whereas our model disentangles it from the action sound.
In addition, we also evaluate the following ablations: “w/o vocoder": we replace the trained HiFi-GAN vocoder with Griffin-Lim; “w/o cond": we remove the audio condition at training time; “w/o cond + denoiser": we use an off-the-shelf model to denoise the target audio  [^3] https://github.com/timsainb/noisereduce; “w/ random test cond": we use random audio from the training set as the condition instead of retrieving audio with the highest AV-Sim score.
@@
Figure 6: Qualitative example. We show the frames of each video followed by the waveform/spectrogram of various baseline methods. Our model generates the most synchronized sounds.
@@ 5.2 Results on Ego4D-Sounds
In this section, we evaluate the ambient-sound joint generation setting with retrieval augmented generation. The results are shown in Tab. 2. Compared to all three baselines, we outperform them on all three metrics by a large margin. While the Retrieval baseline retrieves natural sounds from the training set and has a low FAD score compared to Spec-VQGAN and Diff-Foley, both its AV-Sync accuracy and CLAP scores are very low. Diff-Foley has a higher performance than Spec-VQGAN since it has been trained on this task, but it still largely underperforms our model w/o cond, likely because their video features do not generalize to the egocentric setting well.
@@
We show two qualitative examples in Fig. 6 comparing our model with several baselines and we show that our model synthesizes both more synchronized and more plausible sounds. To fully evaluate our results, it is important to view the supplementary video.
@@
Figure 7: The achieved ambient level and accuracy of the prediction as a function of the input ambient levels. (a): we show the ambient level of our model changes according to the ambient level in the audio condition while the ambient level of “Ours w/o cond" and the original audio stay constant, illustrating the controllability of our model. (b) FAD is low for most input ambient levels unless it goes too extreme (too low or too high), showing our model generates high-quality action sounds even when varying output ambient levels.
@@ 5.3 Ambient Sound Control
By disentangling action sounds from ambient sounds, our model allows taking any given sound as the condition at test time. To examine whether our model truly relies on the audio condition to learn the ambient sound information, we test the model by providing audio conditions of various ambient levels and then calculate the ambient level in the generated audio. The ambient level is defined as the lowest energy of any 0.5s audio segment in a 3s audio.
The results are shown in Fig. 7, where we also plot the ambient levels of “Ours w/o cond" and the original audio. Our model changes the ambient sound level according to the input ambient (shown in Fig. 7(a)) while still synthesizing plausible action sounds (shown in Fig. 7(b)). FAD spikes when the condition ambient is too low or too high, most likely because the generated ambient sound is out of distribution since the original audio always has some ambient sounds.
@@ 5.4 Human Evaluation
To further validate the performance of various models, we conduct a subjective human evaluation. In each survey, we provide 30 questions and each question has 5 videos with the same visuals but different audio samples. For each video, we ask the participant to select the video(s) whose audio 1) is most semantically plausible and temporally synchronized with the video and 2) has the least ambient sounds. We invite 20 participants to complete the survey and compute the average voting for all 30 examples. See the survey interface and guidelines in Supp.
Tab. 3 shows the results. Overall, all learning-based methods generate reasonable action sounds while our model (action-ambient) has the highest score for action-sound quality compared to other methods. Although ours (action-focused) has a slightly lower action-sound score, it has significantly less ambient sound. This is likely because sometimes the low-ambient condition can lead the model to suppress some minor action sounds.
@@ 5.5 Results on EPIC-KITCHENS
To evaluate whether our model generalizes to other datasets, we also test our model on the EPIC-KITCHENS dataset. We first sample 1000 3s clips on EPIC-KITCHENS and then evaluate the retrieval baseline, Diff-Foley, Ours w/o cond, and our full model on these data and then compute the FAD and AV-Sync scores for them.
Results are shown in Tab. 4. Similar to what we observe on Ego4D-Sounds, our model outperforms other models on FAD and AV-Sync by a large margin, showing ours learns better to generate action sounds from visuals, which also transfer to other datasets.
@@
Table 4: Results on Epic-Kichens. GT stands for Ground Truth.
@@ 5.6 Demo on VR Cooking Game
One compelling application of action-to-sound generation is to generate sound effects for games in virtual reality, where simulating complex hand-object interactions is non-trivial. To examine whether our learned model generalizes to VR games, we collect game videos of a cooking VR game “Clash Of Chefs” from YouTube and test our model without fine-tuning. Preliminary results suggest our model can generate synced action sounds (see Fig. 9 and Supp). This suggests promising future in learning action-to-sound models from real-world egocentric videos and applying them to VR games to give a game user an immersive audio-visual experience that dynamically adjusts to their own actions.

### v3

@@ 5 Experiments
    Audio-visual synchronization (AV-Sync) [41]: a binary classification model that classifies whether the video and generated audio streams are synchronized. Following [41], we create negative examples by either shifting audio temporally or sampling audio from a different video clip. Details in Supp.
@@
    REGNET [7]: a video-to-audio model that uses a bottleneck design to generate visually-relevant sounds. We run their trained model on our test set.
  3. 3.
@@
In addition, we provide ablations: “w/o vocoder”: we replace the trained HiFi-GAN vocoder with Griffin-Lim; “w/o cond”: we remove the audio condition at training time; “w/o cond + denoiser”: we use an off-the-shelf model to denoise the target audio  [^4] https://github.com/timsainb/noisereduce; “w/ random test cond”: we use random audio from the training set as the condition instead of retrieving audio with the highest AV-Sim score.
@@
Figure 6: Qualitative examples showing frames followed by the waveform/spectrogram of various baselines. Our model generates the most synchronized sounds.
@@ 5.1 Results on Ego4D-Sounds
First we evaluate the ambient-sound joint generation setting with retrieval augmented generation. Tab. 2 shows the results. Compared to all three baselines, we outperform them on all three metrics by a large margin. While the Retrieval baseline retrieves natural sounds from the training set and has a low FAD score compared to Spec-VQGAN and Diff-Foley, both its AV-Sync accuracy and CLAP scores are very low. Diff-Foley performs better than Spec-VQGAN since it has been trained on this task, but it still largely underperforms our model w/o cond, likely because its video features do not generalize to the egocentric setting well. REGNET performs the worst likely due to its failure to account for strongly present ambient sounds and its assumption for a fixed taxonomy of sounds.
@@
We show two qualitative examples in Fig. 6 comparing our model with several baselines. Our model synthesizes both more synchronized and more plausible sounds. To fully evaluate our results, it is important to view the Supp. video.
@@
Figure 7: The achieved ambient level and accuracy as a function of the input ambient levels. (a) We show the ambient level of our model changes according to the ambient level in the audio condition while the ambient level of “Ours w/o cond” and the original audio stay constant, illustrating the controllability of our model. (b) FAD is low for most input ambient levels unless it goes too extreme (too low or too high), showing our model generates high-quality action sounds even when varying output ambient levels.
@@ 5.2 Ambient Sound Control
By disentangling action sounds from ambient sounds, our model allows taking any given sound as the condition at test time. To examine whether our model truly relies on the audio condition to learn the ambient sound information, we next test the model by providing audio conditions of various ambient levels and then calculate the ambient level in the generated audio. The ambient level is defined as the lowest energy of any 0.5s audio segment in a 3s audio.
Fig. 7 shows the results, where we also plot the ambient levels of “Ours w/o cond” and the original audio. Our model changes the ambient sound level according to the input ambient (shown in Fig. 7(a)) while still synthesizing plausible action sounds (shown in Fig. 7(b)). FAD spikes when the condition ambient is too low or too high, most likely because the generated ambient sound is out of distribution since the original audio always has some ambient sounds.
@@ 5.3 Human Evaluation
To further validate the performance of various models, we conduct a subjective human evaluation. In each survey, we provide 30 questions, each with 5 videos with the same visuals but different audio samples. For each video, we ask the participant to select the video(s) whose audio 1) is most semantically plausible and temporally synchronized with the video and 2) has the least ambient sounds. We invite 20 participants to complete the survey and compute the average voting for all 30 examples. See the survey interface and guidelines in Supp.
Tab. 3 shows the results. All learning-based methods generate reasonable action sounds, yet our model (action-ambient) has the highest score for action-sound quality compared to other methods. Although ours (action-focused) has a slightly lower action-sound score, it has significantly less ambient sound. This is likely because sometimes the low-ambient condition can lead the model to suppress some minor action sounds.
Overall, our model generates both short percussive and longer harmonic action sounds while producing desired ambient sounds controlled by users. The model can fail sometimes in predicting more subtle action sounds, however. See Supp. video for both success and failure examples.
@@ 5.4 Results on EPIC-KITCHENS
To evaluate whether our model generalizes to other datasets, we also test our model on the EPIC-KITCHENS dataset. We first sample 1000 3s clips from EPIC and then evaluate the retrieval baseline, Diff-Foley, Ours w/o cond, and our full model on these data and then compute their FAD and AV-Sync scores.
Tab. 4 shows the results. Similar to what we observe on Ego4D-Sounds, our model outperforms other models by a large margin, showing it better learns to generate action sounds from visuals, even when transferring to another dataset.
@@
Table 4: Results on EPIC-KITCHENS. GT stands for Ground Truth.
@@ 5.5 Demo on VR Cooking Game
One compelling application of action-to-sound generation is to generate sound effects for games in virtual reality, where simulating complex hand-object interactions is non-trivial. To examine whether our learned model generalizes to VR games, we collect game videos of a cooking VR game “Clash Of Chefs” from YouTube and test our model without fine-tuning. Preliminary results suggest our model can generate synced action sounds (see Fig. 9 and Supp). Though there remains much work to do, this suggests a promising future in learning action-to-sound models from real-world egocentric videos and applying them to VR games to give a game user an immersive audio-visual experience that dynamically adjusts to their own actions.
@@ Acknowledgments:
  6. 6.
    Effect of the size of the retrieval pool.
  7. 7.
    Ablation study on video frame rates.
@@ 7.5 Human Evaluation Details
7.6 Effect of the Size of the Retrieval Pool
To understand how the size of the retrieval pool affects the model performance, we evaluate our model with varying pool sizes (10k/1k/100 randomly sampled from the full retrieval pool, and the FAD scores are 1.01/1.02/1.12 (lower is better). We see that the model is fairly robust to the retrieval pool size.
7.7 Ablation Study on Video Frame Rates.
In the paper, we use 16 frames for 3s video (around 5 FPS). We experimented with higher FPS (8 and 10) previously and did not observe significant improvements, likely because most human actions do not have higher frequencies and can be captured with 5 FPS.